OpenStudy (heart_offire):

Write an equation of the line that has a slope of -2/3 and a y-intercept of -2. A) 2x + 3y = -6 B) 2x + 3y = 6 C) 2x - 3y = -6 D) 2x - 3y = 6 @retirEEd

OpenStudy (tkhunny):

Have you considered writing the equation of the line in Slope-Intercept form and then using your best algebra to put it in Standard Form?

OpenStudy (heart_offire):

no because i dont really no how to start

OpenStudy (heart_offire):

@jabez177

OpenStudy (heart_offire):

@Zenmo

OpenStudy (heart_offire):

@mazmaz

OpenStudy (heart_offire):

@pooja195

OpenStudy (tkhunny):

I just told you how to start. Why not try that. Do you know the Slope-Intercept form?

OpenStudy (heart_offire):

is it y=mx+b

OpenStudy (tkhunny):

Right. You are given the slope and intercept in the problem statement. Go!

OpenStudy (heart_offire):

so i then multiply -2/3 and -2

OpenStudy (heart_offire):

because i put it in slope intercept form

OpenStudy (tkhunny):

Why don't you first substitute the slope and intercept into the y = mx+b equation and let's see what you get?

OpenStudy (heart_offire):

y= -2/3x+ -2

OpenStudy (tkhunny):

Good. Now, multiply everything by 3.

OpenStudy (heart_offire):

-2 x 3= -6, -2/3 x 3= -2

OpenStudy (tkhunny):

,,,and y*3 = 3y Put it all together in your new equation.

OpenStudy (heart_offire):

2x+3y=-6

OpenStudy (tkhunny):

Sooo... Are you done?

OpenStudy (heart_offire):

yea but its not one of the options

OpenStudy (photon336):

To find the y intercept all you need to do is plug in x = 0 into the equation. the equation is in the form y = m(x)+b where m = the slope and the y intercept is found by m(0)+b = y But it's not that straightforward because you need to do some manipulation to find your answer. in each of those problems you need to solve for y